The measure of 2 supplementary angles are in the ratio 1:9. Find the degree measure of the smaller angle.

Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12569)

1N+9N=180
10N=180
N=180/10
N=18
1*18=18 DEGREES FOR THE SMALLER ANGLE.
9*18=162 DEGREES FOR THE LARGER ANGLE.
PROOF:
18 + 162 = 180

Answer by Alan3354(31487)

The measure of 2 supplementary angles are in the ratio 1:9. Find the degree measure of the smaller angle.
----------------------
Supplementary angles add to 180º
------------
1a + 9a = 180
a = 18
----
Angles are 18º and 162º